# Quantitative estimates: How well does the discrete Fourier transform approximate the Fourier transform on $\mathbb{R}$?

### v1

@@ 4.1 Convergence rates
$\Big{\|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
(s1) The error $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\|_{L^{2}}$ relates to classical interpolation of $\hat{f}$ from its samples [5, 6]. Since we consider general weights only for $B_{0}=\operatorname*{sinc}$, we carry out the calculations for $B_{0}$. Direct computations and Plancherel’s formula lead to
$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\cdot-k)\Big{\|}^{2}_{L^{2}}$ $=\Big{\|}f-\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}\,\cdot\,\frac{k}{p}}{\bf 1}_{[-\frac{p}{2},\frac{p}{2}]}\Big{\|}^{2}_{L^{2}}$
@@
This provides the required bound
$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$
To treat the case $i=1,2$ of piecewise constant and piecewise linear interpolation for the polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$, we use results from [15, Thm 3 and Lemma 3]. For $\alpha\leq i,i=1,2,$ these imply
$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim p^{-\alpha}\|f\|_{L^{2}_{v_{\alpha}}}\,.$ (31)
The continuous embedding $W(\mathcal{C},\ell^{2}_{v_{i}})\subseteq L^{2}_{v_{i}}$ yields $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\|_{L^{2}}\lesssim p^{-i}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{i}})}$.
@@
The restriction to $\alpha=i$, for $i=1,2$, in Part (b) of Theorem 4.1 stems from the bound (31). For stronger polynomial decay $\alpha$, the splines $B_{1}$ and $B_{2}$ need to be replaced by a suitable function $\phi_{\alpha}$, see [5] for examples. We need at least that $\phi_{\alpha}$ interpolates on $\mathbb{Z}$ and that its integer shifts form a Riesz-sequence in $L^{2}$.
The proof for the interpolating function $\phi=\operatorname*{sinc}$ works for all polynomials weights $v_{\alpha}$ with $\alpha>\frac{1}{2}$. Although the $\operatorname*{sinc}$-function has a bad reputation in numerical analysis due to its poor decay, it works well in the context of Fourier approximation. In particular, since we use only finite sums, the slow decay of $\operatorname*{sinc}$ does not cause any convergence issues. We refer to [21, 22] for $\operatorname*{sinc}$-methods in numerical analysis.
5 Minimal requirements
In this section we investigate under which conditions the error $E^{[n]}_{h}(f)$ tends to zero. Although this question may not be of practical value, it is of fundamental interest to know for which function class the Fourier transform can be successfully approximated by the discrete Fourier transform. We answer this question for the norm $\sqrt{p}\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ representing the pointwise error.
@@ 5 Minimal requirements
Proof 5.2.
Following the general approach outlined in Section 3.1 we decompose the error $\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ into a time and a frequency component. As in (13), we have
$\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\|\hat{f}_{\frac{1}{p},n}-(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\|\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}-\operatorname{\mathcal{F}}\!f_{h,n}\|_{\infty}\,.$ (32)
@@
$\sqrt{p}\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}\,,$ (33)
which again splits the error into a time component and a frequency component.
It remains to obtain suitable decay of samples of $\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f}$ and $f-\mathcal{P}\!_{p}f$. We start with the frequency component $\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$. The observation $|\frac{k}{p}|\leq\frac{1}{2h}$, for $k\in[n]$, and assuming without loss of generality that $h\leq 1$ lead to
$\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$ $=\sup_{k\in[n]}\Big{|}\sum_{l\in\mathbb{Z}\setminus\{0\}}\hat{f}(\tfrac{k}{p}+h^{-1}l)\Big{|}\leq\sum_{|l|\geq\frac{1}{2h}}\sup_{\xi\in[0,1]}|\hat{f}(\xi+l)|\,.$ (34)
@@
Since $f\in W(\mathcal{C},\ell^{1})$, the series tends to zero for $p\rightarrow\infty$. The factor $h\lceil h^{-1}\rceil\leq(1+h)$ does not cause any issues because $h$ is small, after all, we let $h\rightarrow 0$.
Since $\|y\|_{2}\leq\sqrt{n}\,\|y\|_{\infty}$ for $y\in\mathbb{C}^{n}$, Proposition 5.1 also implies the following $\ell^{2}$-error:
$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\,E^{[n]}_{h}(f)\leq\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\frac{\sqrt{n}}{\sqrt{p}}\,\sup_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}\rightarrow 0\,.$
for $f,\hat{f}\in W(\mathcal{C},\ell^{1})$.
In a similar spirit we can ask when the interpolation in (28) with the $B$-splines $B_{1}$ or $B_{2}$ converges to zero.
@@
If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then
$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\,\sup_{\xi\in\mathbb{R}}\Big{|}\hat{f}(\xi)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\Big{|}=0\,,\qquad i=1,2.$ (35)
A similar result under a stronger condition on $f$ and $\hat{f}$ and the choice $p=1/h=\sqrt{n}$ was derived in [16] for the piecewise linear interpolation with $B_{2}$. Proposition 5.3 improves the main result in [16] by extending the approximation result to a larger class of functions. For the rather non-trivial comparison of the condition $f,\hat{f}\in W(\mathcal{C},\ell^{1})$ to the one in [16] we refer to [17, Theorem 2]).
Proof 5.4.
As in the proof of Theorem 4.1 we split the approximation error into discretization, truncation, and approximation via the discrete Fourier transform as follows:
@@
To estimate the error in $(a1)$, we use the fact that the integer shifts of $B_{i}$, $i=1,2$, form a partition of unity, so that $\operatorname*{supp}(B_{i})\subseteq[-1,1]$ leads to
@@
$\leq\Big{|}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}|\hat{f}(\xi)-\hat{f}(\tfrac{k}{p})|B_{i}(p\xi-k)\Big{|}$
$\leq\sup_{k\in\mathbb{Z}}\sup_{\xi\in\mathbb{R}:|\xi-\frac{k}{p}|\leq\frac{1}{p}}\Big{(}\sum_{k\in\mathbb{Z}}B_{i}(p\xi-k)\Big{)}$
$\leq\sup_{\xi,\eta\in\mathbb{R}:|\xi-\eta|<\frac{1}{p}}|\hat{f}(\xi)-f(\eta)|$
@@
For the error in $(a2)$, we use $\|B_{i}\|_{\infty}=1$ and find for all $\xi\in\mathbb{R}$ that
@@
$\leq\|B_{i}\|_{\infty}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\setminus[n]\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}\big{|}\hat{f}(\tfrac{k}{p})\big{|}\leq\sum_{k\in\mathbb{Z}\setminus[n]}\big{|}\hat{f}(\tfrac{k}{p})\big{|}\,.$
@@
In order to estimate $(a3)$, we compute
$\bigg{|}\sum_{k\in[n]}\Big{(}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{)}B_{i}(p\xi-k)\bigg{|}$
@@
were we have used again that the integer translates of the cardinal $B$-splines form a partition of unity. According to Proposition 5.1 this term vanishes when $p\rightarrow\infty$ and $h\rightarrow 0$.
@@
Appendix A Time frequency symmetry of the error functional
We prove the time-frequency symmetry of the error $E^{[n]}_{h}$.
Lemma A.1.
If $p=nh$, then
$E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}}),\qquad f,\hat{f}\in W(\mathcal{C},\ell^{1})\,.$ (36)
Hence, the time frequency decomposition of the error matches the time-frequency symmetry of $E^{[n]}_{h}$.
Proof A.2.
$E^{[n]}_{p^{-1}}(\bar{\hat{f}})$ $=\|(\overline{\widehat{f_{h,n}}})\,\widehat{}-\mathcal{F}\bar{\hat{f}}_{\frac{1}{p},n}\|=\|\bar{f}_{h,n}-\mathcal{F}\bar{\hat{f}}_{\frac{1}{p},n}\|$
=*$=\|\mathcal{F}^{*}\bar{f}_{h,n}-\bar{\hat{f}}_{\frac{1}{p},n}\|=E^{[n]}_{p^{-1}}(\bar{\hat{f}})\,.$
The discrete Fourier transform also satisfies $\overline{\mathcal{F}y}=\mathcal{F}^{*}\bar{y}$, where conjugation is meant entry-wise. Conjugation does not affect the norm, so that we obtain (36).

### v2

@@ 5 Interpolation
$\Big\|\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big\|_{L^{2}}\lesssim\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
(s1) The interpolation error $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\|_{L^{2}}$ relates to classical interpolation of $\hat{f}$ from its samples [6, 7, 28]. We carry out the calculations for $B_{0}=\operatorname*{sinc}$. Direct computations and Plancherel’s formula lead to
$\Big\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\xi-k)\Big\|^{2}_{L^{2}}$ $=\Big\|f-\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}x\frac{k}{p}}{\bf 1}_{[-\frac{p}{2},\frac{p}{2}]}\Big\|^{2}_{L^{2}}$
@@
This provides the bound for the interpolation error as
$\Big\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$
To treat the cases $i=1,2$ of piecewise constant and piecewise linear interpolation for the polynomial weight $v=v_{\alpha}$, we use results from [19, Theorem 3 and Lemma 3]. For $\alpha\leq i,i=1,2,$ these imply
(39) $\Big\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big\|_{L^{2}}\lesssim p^{-\alpha}\|f\|_{L^{2}_{v_{\alpha}}}\,.$
The continuous embedding $W(\mathcal{C},\ell^{2}_{v_{\alpha}})\subseteq L^{2}_{v_{\alpha}}$ yields
$\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\|_{L^{2}}\lesssim p^{-\alpha}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}\,.$
@@
The restriction to $\alpha\leq i$, for $i=1,2$, in the second part of Theorem 5.1 stems from the bound (39). For stronger polynomial decay $\alpha$, the splines $B_{1}$ and $B_{2}$ need to be replaced by a suitable function $\phi_{\alpha}$, see [6] for examples. We need at least that $\phi_{\alpha}$ interpolates on $\mathbb{Z}$ and that its integer shifts form a Riesz-sequence in $L^{2}$.
The proof for the interpolating function $\phi=\operatorname*{sinc}$ holds for all polynomial weights $v_{\alpha}$ with $\alpha>\frac{1}{2}$. The $\operatorname*{sinc}$-function arises naturally in Fourier approximation, but its slow decay poses challenges in numerical analysis. Nevertheless, it has been successfully employed in various numerical methods – see Stenger’s work [27, 28] for comprehensive treatments of $\operatorname*{sinc}$-based techniques.
6 Minimal requirements for convergence
In this section we aim to identify the weakest conditions under which the Fourier transform can be successfully approximated by the discrete Fourier transform. Although this question may not be of immediate practical value, it is a fundamental question of intrinsic mathematical interest.
To identify weak conditions, we now choose the weaker norm
(40) $\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\,$
and ask for sufficient conditions on $f$ and $\hat{f}$ such that (40) converges to 0 when $h\rightarrow 0$ and $p\rightarrow\infty$. It turns out that the same general conditions for the validity of the Poisson summation formula also guarantee the convergence.
@@ 6 Minimal requirements for convergence
The limit above is taken in the sense that $h\to 0$ and $p^{-1}\to 0$ in arbitrary relation to each other. In particular, it implies that $n=\frac{p}{h}\rightarrow\infty$.
Proof 6.2.
Following the general approach outlined in Section 4.1, we decompose the error $\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|=\sqrt{p}\,\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ into a time and a frequency component. Since we are taking limits, we may assume without loss of generality that $h\leq 1$.
As in (4.6), we have
(41) $\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\|\hat{f}_{\frac{1}{p},n}-(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\|\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}-\operatorname{\mathcal{F}}\!f_{h,n}\|_{\infty}\,.$
@@
(42) $\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\leq\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}\,$,
which splits the error into a time component and a frequency component as before.
It remains to obtain suitable decay of samples of $\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f}$ and $f-\mathcal{P}\!_{p}f$. We start with the frequency component $\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$. The observation $|\frac{k}{p}|\leq\frac{1}{2h}$ for $k\in[n]$ and the assumption $h\leq 1$ lead to
(43) $\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$ $=\sup_{k\in[n]}\Big|\sum_{l\in\mathbb{Z}\setminus\{0\}}\hat{f}(\tfrac{k}{p}+h^{-1}l)\Big|\leq\sum_{|l|\geq\frac{1}{2h}}\sup_{\xi\in[0,1]}|\hat{f}(\xi+l)|\,.$
@@
Since $f\in W(\mathcal{C},\ell^{1})$, the series tends to zero for $p\rightarrow\infty$. The factor $h\lceil h^{-1}\rceil\leq(1+h)\leq 2$ does not cause any issues because $h\rightarrow 0$.
In previous sections, we considered the $\ell^{2}$-error $E^{[n]}_{h}(f)$, which is stronger than (40). Since $\|y\|_{2}\leq\sqrt{n}\,\|y\|_{\infty}$ for $y\in\mathbb{C}^{n}$, Proposition 6.1 also implies that for $f,\hat{f}\in W(\mathcal{C},\ell^{1})$,
$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\,E^{[n]}_{h}(f)\leq\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\rightarrow 0\,$,
so that an additional factor $\sqrt{h}$ is needed for convergence.
Parallel to Section 5 we now investigate when the interpolation error of $\hat{f}$ in (36) converges to zero.
@@
If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then for interpolation with step functions $B_{1}$ and with piecewise linear functions $B_{2}$ we obtain
(44) $\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\,\sup_{\xi\in\mathbb{R}}\Big|\hat{f}(\xi)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\Big|=0\,,\qquad i=1,2.$
Proof 6.4.
As in the proof of Theorem 5.1 we split the approximation error into interpolation, truncation, and discretization via the discrete Fourier transform as follows:
@@
To estimate the interpolation error $(a1)$, we use the fact that $B_{i}$ has compact support, $\operatorname*{supp}(B_{i})\subseteq[-1,1]$, and that the integer shifts of $B_{i}$, $i=1,2$, form a partition of unity. This leads to
@@
$\leq\sum_{\begin{subarray}{c}k\in\mathbb{Z}\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}|\hat{f}(\xi)-\hat{f}(\tfrac{k}{p})|\,B_{i}(p\xi-k)$
$\leq\sup_{k\in\mathbb{Z}}\sup_{\xi\in\mathbb{R}:|\xi-\frac{k}{p}|\leq\frac{1}{p}}|\hat{f}(\xi)-\hat{f}(\tfrac{k}{p})|\sum_{k\in\mathbb{Z}}\,B_{i}(p\xi-k)$
$\leq\sup_{\xi,\eta\in\mathbb{R}:|\xi-\eta|\leq\frac{1}{p}}|\hat{f}(\xi)-\hat{f}(\eta)|\,.$
@@
For the truncation error $(a2)$ we use $\|B_{i}\|_{\infty}=1$ and find for all $\xi\in\mathbb{R}$ that
@@
$\leq\|B_{i}\|_{\infty}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\setminus[n]\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}\big|\hat{f}(\tfrac{k}{p})\big|$
$\leq\sum_{k\in\mathbb{Z}\setminus[n]}\big|\hat{f}(\tfrac{k}{p})\big|\,.$
@@
For the discretization error $(a3)$, we estimate
$\hskip-56.9055pt\bigg|\sum_{k\in[n]}\Big(\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big)\,B_{i}(p\xi-k)\bigg|$
@@
where we have used again that the integer translates of the cardinal $B$-splines form a partition of unity. According to Proposition 6.1 this term vanishes when $p\rightarrow\infty$ and $h\rightarrow 0$.
@@
Remark 6.5.
(i) The formulation of Proposition 2.6 in the introduction follows, because the condition $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{1+\epsilon}<\infty$ implies that $f\in W(\mathbb{C},\ell^{1})$.
(ii) Convergence in a stronger norm for the choice $p=1/h=\sqrt{n}$ and under stronger conditions on $f$ and $\hat{f}$ was derived in [20] for the piecewise linear interpolation with $B_{2}$. Proposition 6.3 extends the applicability of the main result in [20] to a larger class of functions. For the rather non-trivial comparison of the condition $f,\hat{f}\in W(\mathcal{C},\ell^{1})$ to the one in [20] we refer to [22, Theorem 2].
(iii) For $B_{0}=\operatorname*{sinc}$ the above proof of Proposition 6.3 breaks down at several steps, because the integer shifts of $\operatorname*{sinc}$ are not summable. We do not know whether Proposition 6.3 remains true in this case.